OpenStudy (anonymous):

The graph of f(x) = 2^(x + 4) shifts six units to the right when it is replaced with the graph of f(x) = 2^(x - k). What is the value of k? 2 4 6 10

OpenStudy (anonymous):

@tkhunny @hartnn @ganeshie8 @phi

ganeshie8 (ganeshie8):

f(x) = 2^(x + 4) to shift the graph of f(x) to right by 6 units : simply replace `x` by `x-6`

OpenStudy (anonymous):

so the answer is 6?

ganeshie8 (ganeshie8):

close, but no. we need to find the value of \(k\), not how many units to shift

OpenStudy (cwrw238):

given graph of f(x) shifting this a units to the right gives f(x - a)

OpenStudy (anonymous):

wait, so how do I find the value of k?

ganeshie8 (ganeshie8):

f(x) = 2^(x + 4) replace `x` by `x-6` and tell me wat do u get ?

OpenStudy (anonymous):

I get f(x) = 2^(x-2)

ganeshie8 (ganeshie8):

Excellent ! compare your result with the given second function : f(x) = 2^(x - k)

ganeshie8 (ganeshie8):

k = ?

OpenStudy (anonymous):

2?

ganeshie8 (ganeshie8):

thats it !

OpenStudy (anonymous):

thanks :)

ganeshie8 (ganeshie8):

yw :)